OpenStudy (anonymous):

anyone please find limit x approach to infinity [(x+2)/(x+1)]^x ? thanks

OpenStudy (shamil98):

read this: http://www.mathsisfun.com/calculus/limits-infinity.html gives a nice detailed explanation on this stuff.

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x \to \infty }\left( \frac{x+2}{x+1}\right)^x\) \(\large \lim \limits_{x \to \infty }e^ {\ln \left( \frac{x+2}{x+1}\right)^x}\) \(\large e^ {\lim \limits_{x \to \infty } \ln \left( \frac{x+2}{x+1}\right)^x}\) \(\large e^ {\lim \limits_{x \to \infty } x\ln \left( \frac{x+2}{x+1}\right)}\)

OpenStudy (anonymous):

okay..thank you very much... :-*

ganeshie8 (ganeshie8):

np :) thats not all.... still so much to do ;)

OpenStudy (anonymous):

yaa okay.. i'll try find the answer.. ;)

ganeshie8 (ganeshie8):

one easy way is to use the known limit : lim x-> inf (x+1/x)^x = e

ganeshie8 (ganeshie8):

substitute x+1 =t, and convert the given limit to that form and conclude

OpenStudy (raden):

((x+1+1)/(x+1))^x = (1 + 1/(x+1))^x+1-1 = (1 + 1/(x+1))^x+1 * (1 + 1/(x+1))^-1 lim x->~ (1 + 1/(x+1))^x+1 * (1 + 1/(x+1))^-1 = limx->~ (1 + 1/(x+1))^x+1 * limx->~ (1 + 1/(x+1))^-1 now you can identify that what's the answer then ...

OpenStudy (anonymous):

okay..thank you yaa... ;) ;)

OpenStudy (raden):

sepertinya anda orang melayu, ya :v are you from malaysian or indonesian ??

OpenStudy (anonymous):

hahaha iyaaaa..malaysia..makasih yaa..

OpenStudy (raden):

okay, kita bertetangga :) saye dri indonesia

OpenStudy (anonymous):

oh..indonesia.. met kenalan yaa.. makasih sekali lagi buat jawapannya..

OpenStudy (raden):

jadi, apa jawabanmu adik manis, adik comel ?

OpenStudy (anonymous):

ini soalan dari adik saya..dia sedang jawab.. ;)